Q.a rubber ball dropped from a height of 20m(g=10m/s2) falls on the roof of an elevator going up at 2m/s.if the collision of the ball with the elevatior is elastic,the rebounds with a velocity of?


ANS:-- 24m/s

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

5t2+2t=20, 5t2 +2t-20=0,t=9/5sec,v=gt =18m/s, relative velocity at collision = 18+2=20m/s, therefore velocity of rebound =20+2=22m/s